OpenStudy (anonymous):

Distinct four-letter sequences are formed by picking 4 letter tiles from a bag containing 11 different alphabet tiles. Note that the order in which the letters are picked matters. The probability of getting a particular four-letter sequence is ___________

OpenStudy (anonymous):

Do you know whether or not the tiles are put back into the bag once they are drawn?

OpenStudy (anonymous):

I dont think they are, because the question doesnt say theyre put back

OpenStudy (anonymous):

nope im pretty sure theyre not put back

OpenStudy (anonymous):

Great! So lets start with the first tile, and work our way up to the 4 tile sequence. The first time you draw a tile, there are 11 tiles in the bag. So, assuming that they are random and each tile is unique, the probability of drawing any particular tile is 1 in 11 or 1/11 or 9.0909090909090...% But now that tile is no longer in the bag and we only have 10 tiles left. So now the probability of drawing the second tile is 1 in 10 or 10% By that reasoning, the chance that the 3rd tile is a specific tile is 1/9 and the chance that the 4th tile is a specific tile is 1/8. The overall probability is the product of these probabilities because first you need the 1 in 11 to happen, then the 1 in 10, then the 1 in 9, then the 1 in 8. So the probabilities compound and your overall probability is: p = 1/11 * 1/10 * 1/9 * 1/8 I'll let you do the math to simplify it!

OpenStudy (anonymous):

\[\frac{ 1 }{ 11 }*\frac{ 1 }{ 10 }*\frac{ 1 }{ 9 }*\frac{ 1 }{ 8 }=\frac{ 1 }{ 7920 } \] so that's my answer?